Fermi National Laboratory
 

Operation Summary for April 20, 2011 - April 22, 2011

 

Wednesday April 20

The day shift began with the Tevatron (TEV) off for accesses, with Operations monitoring the Recycler stash (331.3E10), with the Antiproton Source (Pbar) stacking (8.5E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTFB T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) in access.

During the day shift, CDF and D0 came out of access at 9:20 AM; operators started the TEV turn-around. Operations disabled beam to T-978 and put the TEV into HEP shot setup at 12:45 PM. Operations established store 8680 at 2 PM with an average initial luminosity (AIL) of 353E30. Operations resumed sending beam to T-978. FESS personnel brought the MI-20 LCW pump back into service. RF technicians reconnected a Booster RF station (BRF4) to the Booster East Anode Power Supply. MTA - MuCool personnel began beam studies at 2:45 PM.

During the evening shift, Operations disabled beam to T-978 at 6 PM. MTA personnel completed their studies at 7:10 PM.

Thursday April 21

The midnight shift began with Operations monitoring store 8680 (103.63E30) and the stash (278.5E10), with Pbar stacking (20.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-978 at 5:05 AM. Operations terminated store 8680 at 5:40 AM and put the TEV into shot setup at 5:50 AM, but abandoned the setup at 6 AM when a low beta quadrupole power supply (D0Q6) tripped off on a ground fault indication. EE Support personnel came in and confirmed an excess current draw in the supply.

During the day shift, Operations put the TEV into shot setup at 10:35 AM and established store 8683 at 12:45 PM with an AIL of 370E30. Controls personnel replaced a ramp control card for several NuMI corrector magnet power supplies, which caused about an hour of downtime.

During the evening shift, Operations disabled beam to T-978 around 6 PM. MTA personnel established high intensity beam to their target. A Linac RF station (LRF5) power amplifier (PA) failed; Linac personnel came in and started preparations to replace the PA.

Friday April 22

The midnight shift began with Operations monitoring store 8683 (101.33E30), the stash (288.1E10), and the Pbar stack (8.0E10), with Linac off due to LRF5 repairs, and with NuMI and MiniBooNE waiting on beam.

During the midnight shift, Linac personnel reported that they had replaced the PA for LRF5. Linac resumed sending beam at 12:55 AM. Operations established beam to T-978 at 4:15 AM. Operations terminated store 8683 at 5:05 AM, put the TEV into shot setup, and established store 8684 at 6:22 AM with an AIL of 361E30.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow cryo system personnel to begin replacing the flywheel on the TEV sector A4 wet engine after store 8684 is terminated at 10 AM. He will also allow MI, Pbar, and Booster personnel to conduct studies.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/22/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory